OpenStudy (anonymous):

how is factoring used to solve quadratic equations. Need to demonstrate the process with an example to show.

OpenStudy (anonymous):

hi nikki

OpenStudy (anonymous):

factoring is used every day

OpenStudy (anonymous):

Hi

OpenStudy (anonymous):

wazzup?

OpenStudy (anonymous):

you asked the same question twice

OpenStudy (anonymous):

Just having trouble with this class

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I know...for some reason, it happened

OpenStudy (anonymous):

well if you throw a baseball with initial velocity v0, and initial height y0, you will need to factor to find when it hits the ground

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

the height of the ball is y = -1/2 gt^2 + vo*t + y0 , where g is your acceleration due to gravity (makes the ball come back to ground)

OpenStudy (amistre64):

you wanna know about factoring right?

OpenStudy (anonymous):

you asked where do we factor? when we solve trajectory problems you dummy

OpenStudy (anonymous):

ask me huh again, i dare you

OpenStudy (anonymous):

wow...that is mean

OpenStudy (anonymous):

im not getting paid after all

OpenStudy (anonymous):

amistre, super hero? wow

OpenStudy (anonymous):

how is factoring used to solve quadratic equations. Need to demonstrate the process with an example to show. Is what I asked...I just do not understand what they want from me...

OpenStudy (amistre64):

solving for any "roots" or "zeros" of an equation, what we are doing is finding all the places that the graph hits the x axis because y = 0 there.

OpenStudy (anonymous):

amistre, its useless

OpenStudy (anonymous):

he will just say "huh"

OpenStudy (amistre64):

with quadratics, the line of the graph bends in such a manner that it can cross the x axis twis, so there may be at most 2 zeros to solve.

OpenStudy (anonymous):

hmmm, i might have misread the question after all, my bad

OpenStudy (anonymous):

amistre, can you solve the surface area of a torus, using integral 2pi y dS?

OpenStudy (anonymous):

amistre, check the discriminant b^2 - 4ac, there are three cases

OpenStudy (amistre64):

for example, can you see the place that this quadratic curve touches the x axis in two spots?

OpenStudy (anonymous):

amistre, youre like god here. is youre mathematical knowledge proportionate to your score?

OpenStudy (anonymous):

I appreciate all the help...I did not understand what they wanted...I'm sorry if it made anyone mad...

OpenStudy (amistre64):

im just an idiot in disguise ;)

OpenStudy (anonymous):

nikki, no one is mad :)

OpenStudy (anonymous):

sometimes im mean, but im not mad :)

OpenStudy (anonymous):

I am not a guy either...lol

OpenStudy (anonymous):

oh sorry

OpenStudy (anonymous):

wait, so youre a goddess :)

OpenStudy (anonymous):

Its ok...sometimes I just do not get this stuff...I want to...it just does not happen...Not a math person...

OpenStudy (anonymous):

amistre, surface area integrals. driving me nuts!!!

OpenStudy (anonymous):

A mother of 2

OpenStudy (anonymous):

i have it here on twiddla http://www.twiddla.com/521818

OpenStudy (amistre64):

when we factor the quadratic equation we get if "two" component parts that multiplied to get it: for example: x^2 +6x +8 is a quadratic expression right?

OpenStudy (anonymous):

i think the student is lost in the math jargon. it happens

OpenStudy (anonymous):

I think so

OpenStudy (amistre64):

surface area integrals tend to be of the "ln(sec+tan)" variety right?

OpenStudy (anonymous):

I HATE MATH!!!!!!!! Just sayin!

OpenStudy (anonymous):

hmmm, not really , could be, check this link http://www.twiddla.com/521818 i posted the problem

OpenStudy (anonymous):

nikki, you dont hate math. you just want things to be logical. i understand

OpenStudy (amistre64):

Nikki: when we "factor" the quadratic we get this: (x+2)(x+4) = 0 we make it equal to zero because we want to know what x is when y = 0

OpenStudy (amistre64):

cantor; ill check it out in abit ;)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

nikki, whats a quadratic equation> what it means to "solve quadratic equation"

OpenStudy (anonymous):

Oh yes...logical would be great:) I am one that needs to get it right now, I am a perfectionist and this is not helping anything, throws me off...very frustrating!

OpenStudy (anonymous):

ok , slow down. ok lets get the jargon straight

OpenStudy (anonymous):

a quadratic equation is generally y = ax^2 + bx + c

OpenStudy (anonymous):

ill show how factoring helps later to solve, but what does it mean to solve a quadratic equation. this is more jargon or math lingo

OpenStudy (amistre64):

Nikki: we know that any number when multiplied by "0" = 0 right? for example: 3(0) = 0 34(0) = 0 r2d2(0)= 0 you with me?

OpenStudy (anonymous):

sure...

OpenStudy (anonymous):

amistre, want to chime in. what it means to "solve quadratic equation"

OpenStudy (amistre64):

cantor, to find their roots :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the roots are the x solutions such that ax^2 + bx + c = 0

OpenStudy (amistre64):

Nikki; then when we have the factored form of a quadratic, we know that one or the other has to equal zero in order for function to equal zero. tell me: (x+4)(0) = 0 ? does (0)(x+2) = 0?

OpenStudy (anonymous):

ok, i got lost there for a second. ok a quadratic equation in x is ax^2 + bx + c = some number, correct? a quadratic function i think would be y = ax^2 + bx + c , because now you have y is a function of x

OpenStudy (anonymous):

ok...I think I got it....hopefully:) Thanks for all your help!

OpenStudy (anonymous):

amistre, ill politely bow out of here, but please check my twiddla

OpenStudy (amistre64):

ill be there shortly, and maybe I can even be helpful ;)

OpenStudy (anonymous):

oh i didnt state the question, one sec

OpenStudy (anonymous):

surface area of a torus. it is generated by circle with radius r , off center. Suppose the circle is centered at (x - R)^2 + y^2 = r^2, and R> r. so the distance from (0,0) to the center of the circle is (R,0) which has radius r. Ok the integral i got is int 2pi y sqrt [ 1+ (dy/dx)^2] = int 2pi sqrt ( r^2 - ( x-R)^2) [ 1 + (-2(x-R)/(2sqrt( r^2 - ( x-R)^2) ^2 ]

OpenStudy (anonymous):

Is this to me?

OpenStudy (amistre64):

Nikki; the factored form allows us to "see" into the quadratic, and know quite easily what the solutions are. (x+4)(x+2) = 0 x = -4 or x = -2 because both these solutions makes one or the other a (0) right?

OpenStudy (anonymous):

yeah

OpenStudy (amistre64):

ack!!...a torus surface..... lol

OpenStudy (amistre64):

Nikki: then that is all ther eis to it, I just hope you can make sense of it between the posting lol

OpenStudy (anonymous):

I am so confused...not sure if it was all for me or for you guys?lol

OpenStudy (anonymous):

well i have two ways to solve it, one way works, the other doesnt, int 2pi y ( 1 + dy/dx^2) dx , that doesnt work, but int 2pi x ( 1 + dx/dy^2) dy that works. not sure why

OpenStudy (anonymous):

nikki, ok i will start over , want to do this? ARE YOU READY!!!

OpenStudy (anonymous):

YES! Sorry...

OpenStudy (anonymous):

oh no im gonna miss jay leno

OpenStudy (anonymous):

ok , uno momento

OpenStudy (anonymous):

I am glad I have DVR

OpenStudy (anonymous):

what do you watch?

OpenStudy (anonymous):

a lot of shows, I really like greys and things like that...so I record them so I do not have to watch all the commercials

OpenStudy (amistre64):

I see the equations..... wheres the torus?

OpenStudy (anonymous):

ok whats a quadratic equation In mathematics, a quadratic equation is a polynomial equation of the second degree. The general form is ax^2 + bx + c = 0

myininaya (myininaya):

cantorset, i wrote log3(n)<n as n<3^n for the log proof sorry for interrupting later

OpenStudy (anonymous):

Are you still there?

OpenStudy (anonymous):

hey sorry

OpenStudy (anonymous):

Its ok....I have to figure out an example for this, and I guess I am not getting this as much as I wanted to be or thought...

OpenStudy (anonymous):

give me 5 minutes

OpenStudy (anonymous):

ok....I appreciate your help...I really do.

OpenStudy (anonymous):

I will go check on my daughters

OpenStudy (anonymous):

ok I'm back.....

OpenStudy (amistre64):

welcome back :)

OpenStudy (anonymous):

Thanks! I am trying to figure out how some get this math stuff, and then there is me...who has no idea and thinks I get it and then does not...